Two hikers are 12 miles apart and walking towards each other. They meet in two hours. Find the rate of each hiker if one hiker walks 1.8 mph faster than the other.
Answer by richwmiller(17219)

12 =2*(x)+2*(x+1.8)
x = 2.1
x+1.8=2.1+1.8=3.9